# FedEx: Expediting Federated Learning over Heterogeneous Mobile Devices by Overlapping and Participant Selection

### v1

@@ 5.5. Impacts of Different Staleness Ceilings
From the results, we find that the staleness ceilings trade-off double-edges of overlapping/continuous computing (i.e., training speedup and model drift) across various tasks, which validates our analysis in Sec. 3.5. While overlapping helps to accelerate FL training, it suffers from model drift issues in heterogeneous environments, which may conversely slow down the FL convergence and potentially diminish the overlapping speedup benefits. Specifically, if the value of staleness ceiling is too small, say $0.5K$, it restricts model drift but limits overlapping acceleration, which results in slow convergence. Meanwhile, if the value of staleness ceiling is too big, e.g., $30$ ($3K$), it makes model drift dominant, which also results in slow convergence. Aside from two extreme ends, an appropriate selection of staleness ceilings will help overlapping speedup benefits outweigh the convergence slowdown due to model drift, and effectively reduce the FL training latency. Besides, the good choice of staleness ceiling is learning task specific. For example, as shown in Fig. 7, for the CNN@MNIST and CNN@HAR, $U=10$ (around $K$) has the best performance. For Squeezenet@CIFAR10 and LSTM@Shakespeare, a relatively large staleness ceiling (i.e., $U=20$, around $2K$) would be favored.
@@
Figure 8. FedEx sensitivity with different $\alpha$ values.
5.6. FedEx Sensitivity Analysis
FedEx uses the scaling coefficient $\alpha$ to balance overlapping latency utility and statistical utility, where it adaptively prioritizes high overlapping latency utility participating devices, and penalizes the utility of stragglers in PS. Figure 4 and Table 2 show that FedEx ($\alpha=2$) outperforms its counterparts, and Fig. 8 shows that FedEx achieves similar performance across all non-zero $\alpha$ values ($\alpha=1,2,5$ or $10$). The results validate that FedEx can orchestrate its components to automatically navigate for the desired FL performance and training performance is not that sensitive to $\alpha$ values. When setting $\alpha=0$, the FL server solely relies on statistical utility to select participating devices, which overlooks the time-saving benefits brought by overlapping yet amplifies its impacts of model drifts, so it has even worse performance than FedAvg.
@@ 5.6. FedEx Sensitivity Analysis
Figure 9. Impacts of data heterogeneity.
5.7. Impacts of Data Heterogeneity
We further evaluated the performance of FedEx under various non-i.i.d. levels of training data samples. Figure 9 shows the testing accuracy and corresponding speedups under non-i.i.d. levels of $\lambda=0$, $0.5$ and $1$, respectively, where the latency of FedAvg is the benchmark. Compared with other designs, FedEx always has the most speedup to reach the target accuracy, especially at $\lambda=0.5$. In particular, for the i.i.d. data distribution case (i.e., $\lambda=0$) on CNN@MNIST, we find that FedEx is better than Oort or DGA in terms of the training latency to reach the target accuracy. The potential reason is that the computing and wireless heterogeneity still exist among selected mobile devices, even though the data is i.i.d.. FedEx considers the latency reduction both by overlapping and by addressing straggler issues per round, and works well in such heterogeneous environments. Besides, for the non-i.i.d. cases on CNN@MNIST, DGA has the worst performance (below 88% lower bound of testing accuracy shown in Fig. 9) or even fail FL convergence because of deficiencies discussed in Sec. 2.2. As for FedEx, Oort and FedAvg, we observe the same performance trend across various $\lambda$ values for Squeezenet@CIFAR10, as shown in Fig. 9(b). For DGA, its performance is not included in Fig. 9(b), since DGA’s testing accuracy remains below $50\%$ across different $\lambda$ values in heterogeneous environments. In particular, for the i.i.d. case, the convergence speed of DGA is even slower than that of FedAvg. This is because DGA requires all devices to participate in each round, which involves the straggler every time in the heterogeneous FL setting. FedAvg mitigates this issue by randomly sampling devices for participation in each FL training round.
Table 5. Impact of system heterogeneity.
(H: High-end, Xiaomi 12S; M: Middle-end, Honor 70; L: Low-end, NVIDIA TX2)
@@ 5.7. Impacts of Data Heterogeneity
5.8. Impacts of System Heterogeneity
@@
We also explore the performance of FedEx under different types of heterogeneity, as shown in Table 5. We select three devices with significantly different computational and communication capabilities for emulation experiments: Xiaomi 12S as High-end (H), Honor 70 as Middle-end (M), and NVIDIA TX2 as Low-end (L). Keeping the total number of devices at 100, we set up three device distribution schemes: 70H/20M/10L (indicating 70 High-end, 20 Middle-end, and 10 Low-end devices); 30H/30M/40L; and 20H/30M/50L. We compare the overall latency (OL) to reach the target accuracy in FL training under these three settings using LSTM@Shakespe-are. The experimental results show that as the number of Low-end devices increases, the OL increases for all methods. In all types of system heterogeneity, FedEx significantly outperforms Oort and FedAvg. Since the computation latency of the Low-end TX2 is much greater than its communication latency, the benefit from overlap training decreases as the proportion of such devices increases, resulting in a reduced improvement from FedEx.
6. Conclusion

### v3

@@ V-E Impacts of Different Staleness Ceilings
From the results, we find that the staleness ceilings trade-off double-edges of overlapping/continuous computing (i.e., training speedup and model drift) across various tasks, which validates our analysis in Sec. III-E. While overlapping helps to accelerate FL training, it suffers from model drift issues in heterogeneous environments, which may conversely slow down the FL convergence and potentially diminish the overlapping speedup benefits. Specifically, if the value of staleness ceiling is too small, say $0.5K$, it restricts model drift but limits overlapping acceleration, which results in slow convergence. Meanwhile, if the value of staleness ceiling is too big, e.g., $30$ ($3K$), it makes model drift dominant, which also results in slow convergence. Aside from two extreme ends, an appropriate selection of staleness ceilings will help overlapping speedup benefits outweigh the convergence slowdown due to model drift, and effectively reduce the FL training latency. Besides, the good choice of staleness ceiling is learning task specific. For example, as shown in Fig. 7, for the CNN@MNIST and CNN@HAR, $U=10$ (around $K$) has the best performance. For Squeezenet@CIFAR10 and LSTM@Shakespeare, a relatively large staleness ceiling (i.e., $U=20$, around $2K$) would be favored.
@@
Figure 8: FedEx sensitivity with different $\alpha$ values.
V-F FedEx Sensitivity Analysis
FedEx uses the scaling coefficient $\alpha$ to balance overlapping latency utility and statistical utility, where it adaptively prioritizes high overlapping latency utility participating devices, and penalizes the utility of stragglers in PS. Figure 4 and Table II show that FedEx ($\alpha=2$) outperforms its counterparts, and Fig. 8 shows that FedEx achieves similar performance across all non-zero $\alpha$ values ($\alpha=1,2,5$ or $10$). The results validate that FedEx can orchestrate its components to automatically navigate for the desired FL performance and training performance is not that sensitive to $\alpha$ values. When setting $\alpha=0$, the FL server solely relies on statistical utility to select participating devices, which overlooks the time-saving benefits brought by overlapping yet amplifies its impacts of model drifts, so it has even worse performance than FedAvg.
@@ V-F FedEx Sensitivity Analysis
Figure 9: Impacts of data heterogeneity.
V-G Impacts of Data Heterogeneity
We further evaluated the performance of FedEx under various non-i.i.d. levels of training data samples. Figure 9 shows the testing accuracy and corresponding speedups under non-i.i.d. levels of $\lambda=0$, $0.5$ and $1$, respectively, where the latency of FedAvg is the benchmark. Compared with other designs, FedEx always has the most speedup to reach the target accuracy, especially at $\lambda=0.5$. In particular, for the i.i.d. data distribution case (i.e., $\lambda=0$) on CNN@MNIST, we find that FedEx is better than Oort or DGA in terms of the training latency to reach the target accuracy. The potential reason is that the computing and wireless heterogeneity still exist among selected mobile devices, even though the data is i.i.d.. FedEx considers the latency reduction both by overlapping and by addressing straggler issues per round, and works well in such heterogeneous environments. Besides, for the non-i.i.d. cases on CNN@MNIST, DGA has the worst performance (below 88% lower bound of testing accuracy shown in Fig. 9) or even fail FL convergence because of deficiencies discussed in Sec. II-B. As for FedEx, Oort and FedAvg, we observe the same performance trend across various $\lambda$ values for Squeezenet@CIFAR10, as shown in Fig. 9(b). For DGA, its performance is not included in Fig. 9(b), since DGA’s testing accuracy remains below $50\%$ across different $\lambda$ values in heterogeneous environments. In particular, for the i.i.d. case, the convergence speed of DGA is even slower than that of FedAvg. This is because DGA requires all devices to participate in each round, which involves the straggler every time in the heterogeneous FL setting. FedAvg mitigates this issue by randomly sampling devices for participation in each FL training round.
TABLE V: Impact of system heterogeneity. (H: High-end, Xiaomi 12S; M: Middle-end, Honor 70; L: Low-end, NVIDIA TX2)
@@ V-G Impacts of Data Heterogeneity
V-H Impacts of System Heterogeneity
@@
We also explore the performance of FedEx under different types of heterogeneity, as shown in Table V. We select three devices with significantly different computational and communication capabilities for emulation experiments: Xiaomi 12S as High-end (H), Honor 70 as Middle-end (M), and NVIDIA TX2 as Low-end (L). Keeping the total number of devices at 100, we set up three device distribution schemes: 70H/20M/10L (indicating 70 High-end, 20 Middle-end, and 10 Low-end devices); 30H/30M/40L; and 20H/30M/50L. We compare the overall latency (OL) to reach the target accuracy in FL training under these three settings using LSTM@Shakespe-are. The experimental results show that as the number of Low-end devices increases, the OL increases for all methods. In all types of system heterogeneity, FedEx significantly outperforms Oort and FedAvg. Since the computation latency of the Low-end TX2 is much greater than its communication latency, the benefit from overlap training decreases as the proportion of such devices increases, resulting in a reduced improvement from FedEx.
VI Conclusion